Officials estimate that 320,000 Boston-area party-goers attended the 1995 Independence Day celebration on the banks of the Charles River. They also estimate that the party-goers left behind 40 tons of garbage. Given that a ton
equals 2,000 pounds, how many pounds of garbage did the average party-goer
leave behind?

Answer by jorel1380(3719)

40x2000/320000=80000/320000=1/4 lb. garbage per person..